A 8.0-V battery is connected in series with a 45-mHinductor, a 150-Ω resistor, and an open switch.

Part A

What is the current in the circuit 0.110 ms after the switch is closed?

Express your answer using two significant figures.

I

=

  mA  

Part B

How much energy is stored in the inductor at this time?

Express your answer using two significant figures.

U

=

  μJ  
Answer #1

a) I(t) = (V / R) * (1 - e-tR/L)

= (8.0 / 150) * (1 - e-0.110 * 10-3 * 150/45*10-3)

current = 0.0164 A = 16.4 mA

b) energy = 1/2 L I2

= 1/2 * 45 * 10-3 * 0.01642

energy stored = 6.03 muJ